OpenStudy (anonymous):

Find the most general function M(x,y) for the equation: M(x,y)dx +(x sec^2y- x/y)dy=0 so that the equation is exact

OpenStudy (anonymous):

is M(x,y) =tany -lny

OpenStudy (anonymous):

For the equation 2 b exact, dM/dy must be equal to dN/dx. So, my 1st try is find dN/dx where\[N = ( x \sec ^{2}y -\frac{ x }{ y})\] therefore \[\frac{ dN }{ dx }= \sec ^{2}y -\frac{ 1 }{ y }\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Hw to get that ans? Can u pls explain further?

OpenStudy (anonymous):

just integrate both sides

OpenStudy (anonymous):

So the next step is integrating dN/dx?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and hence N=tany -lny +c

OpenStudy (anonymous):

but it should be dN/dy